# Personalization Toolkit: Training Free Personalization of Large Vision Language Models

### v2

@@ 4.2 Visual-Question Answering
In Table 3, we evaluate our method’s ability to answer questions about personalized objects using Yo’LLaVA [25] and This-Is-My-Img benchmarks. For completeness we also compare PeKit to MyVLM’s on the MyVLM dataset using CLIPScore [14] and Personalization Recall.
@@
On the MyVLM dataset, PeKit achieves a higher performance than MyVLM method in generaing captions aligned with the images (CLIPScore) and incorporating the correct personalized concept names into its captions (Personalization Recall). On the Yo’LLaVA dataset our PeKit outperforms Yo’LLaVA without requiring training, special tokens, or modifications to the original LVLM.
Concerning This-Is-My-Img, all methods’ VQA performance experience a drop compared to the results on the Yo’LLaVA dataset, indicating the difficulty of our benchmark and room for future work. Due to this challenging nature and objects being sometimes in the background, we observe that Yo’LLaVA fails to improve over base LLaVA model. To the contrary, our PeKit achieves significant improvement of ($\sim 5\%$). This highlights our method strong ability in both visual recognition and question answering.
Figure 4: Qualitative Comparison to Yo’LLaVA: Yo’LLaVA’s prompt template requires specifying the personalized object’s identifier in the query (first row), limiting generalization since users must already know which objects are in the image. Using image-level embeddings can also cause confusion between similar objects (e.g., Alex vs. Alex’s bag). Adjusting the LLM’s head weights further harms captioning quality. PeKit achieves better captioning quality without any training. In each example, the “Query” shows Yo’LLaVA’s prompt with the concept identifier replaced by the object’s name and an added system prompt. PeKit uses a different template, detailed in the Appendix.
4.3 Ablations and Analysis
Choice of $F_{\text{ext}}$. Our method employs an open-vocabulary approach to localize personalized objects in reference images. In Table 1, we ablate the choice of an open-world object detector, G-DINO [24], compared to the open-world semantic segmentation model, G-SAM [30], for $F_{\text{ext}}$ in Eq. 1. The results show that our method outperforms previous methods in both cases. However, the more precise segmentation model, which extracts only patches of the object of interest, achieves the best accuracy on average. This improvement is more pronounced on the MyVLM dataset, where objects are less centered in the training views compared to the Yo’LLaVA dataset.
Other LVLMs. The modularity of our method allows for straightforward plug-and-play integration with any LVLM without tuning. We pair PeKit with InternVL2-26B [7] , a state-of-the-art LVLM. Table 3 shows VQA accuracy on the Yo’LLaVA and This-Is-My-Img datasets for PeKit with LLaVA [22] and InternVL [7]. With both backbones, our approach achieves SOTA results. The improved quality of InternVL [7] boosts VQA accuracy on the Yo’LLaVA dataset and achieves a significant $7\%$ boost on This-Is-My-Img.
Qualitative VQA Results Figure 4 shows examples of our PeKit compared to the Yo’LLaVA for VQA and personalized captioning tasks on This-is-my-Img dataset. As seen on the first row, Yo’LLaVA’s prompt template requires the query to include the target object’s concept identifier, making it unsuitable for general captioning tasks and tailored for Visual Question Answering. Besides, since Yo’LLaVA operates on image-level embeddings of reference views, it needs clutter-free object-centered reference views of the personalized objects. As seen on the second row, performance can decline if the personalized object is not in the foreground or if there are other objects/people interacting with the personalized object in the reference views. Besides, fine-tuning the last layer of the language model reduces the LVLM’s captioning capabilities for Yo’LLaVA.
In contrast, PeKit accurately identifies objects and uses visual prompting to guide LLaVA in answering the questions of ambiguous situations where multiple object are present in the image (e.g, Nikki’s car).
@@ 5 Discussion and Conclusion
In this work, we leverage vision foundation models combined with retrieval-augmented generation and visual prompting to build a training-free plug-and-play toolkit for the task of LVLM personalization called PeKit. We first show that our approach can surpass existing training-based methods while not requiring any tuning and not relying on any additional images other than those used to introduce the personalized concept. We further present a challenging benchmark that mimics more realistic scenarios and increases the complexity of the visual recognition challenge. While our approach shows robustness, there is room for improvement in future works. We believe that our PeKit sets a strong training-free baseline for the LVLM personalization task.

### v3

@@ 4.4 Visual-Question Answering
In Tab. 3, we assess PeKit performance in answering questions about personalized objects across various scenarios. We evaluate our approach using the Yo’LLaVA [20] and This-Is-My-Img benchmarks. For a comprehensive comparison, we also evaluate PeKit against MyVLM on the MyVLM dataset using CLIPScore [10] and Personalization Recall.
@@
Our method consistently outperforms baseline models on both single- and multi-concept VQA tasks. On the MyVLM dataset, PeKit surpasses the MyVLM method in generating image-aligned captions (as measured by CLIPScore) and in correctly incorporating personalized concept names (as measured by Personalization Recall). On the Yo’LLaVA dataset, PeKit outperforms Yo’LLaVA despite requiring no additional training, special tokens, or modifications to the original LVLM.
On the This-Is-My-Img benchmark, Yo’LLaVA fails to improve over the base LLaVA model, whereas PeKit achieves a notable performance gain of approximately 5%, demonstrating strong capabilities in both visual recognition and question answering. Additionally, PeKit generalizes well to the video domain (3.5), unlike prior approaches, which are not readily extendable to video. In this setting, our method improves the original model’s performance by 12% and continue to show gains when paired with a more capable instruction-following backbone.
4.5 Qualitative Results
Fig. 3 presents examples of our PeKit model performing various tasks on the proposed benchmark, This-Is-My-Img. In the left column, PeKit successfully identifies personalized single-concept objects, even in ambiguous scenarios involving multiple instances from the same object category. In the middle column, it demonstrates the ability to answer questions involving multi-concept pairs. Lastly, without any modifications, the model can be applied to video inputs by visually prompting the model with detected personalized objects in the video. Additional qualitative comparisons with the base LLaVA model, MyVLM, and Yo’LLaVA are provided in Appendices E, F, and G.
4.6 Ablation and Analysis
A comprehensive evaluation of our method is provided in Appendix.A, where we analyze robustness under various settings. Specifically, we examine the effects of altering the backbone encoder $F_{emb}$, the number of reference images per object (N), the number of personalized objects (P) in the dataset, the detection threshold ($\tau$), semantic category ($k_{p}$) and the name ($n_{p}$) of the personalized objects.
@@ 5 Discussion and Conclusion
In this work, we introduced PeKit, a training-free, plug-and-play toolkit for LVLM personalization that combines vision foundation models with retrieval-augmented generation and visual prompting. Our approach outperforms existing training-based methods without requiring any fine-tuning or additional data beyond the personalized concept inputs. To evaluate our method, we proposed a challenging new benchmark that reflects more realistic scenarios involving multi-concept and video personalization, significantly increasing the complexity of the visual recognition task. PeKit demonstrates strong robustness and scalability across both multi-concept and video settings. Despite its strengths, our benchmark reveals that there is still room for improvement, particularly in handling complex temporal and multi-concept reasoning. We believe that PeKit establishes a strong training-free baseline for future research in LVLM personalization.
Appendix A Ablation
In this section we ablate various aspects of our personalization method primarily using the Yo’LLaVA dataset as it serves as a well-established benchmark.
A.1 Retrieval Module Backbone
We present an ablation study of the image encoder $F_{emb}$, which serves as the backbone of the Retrieval Module and is responsible for extracting object features and performing matching. We observe a noticeable gap between CLIP and DINO, indicating that DINO’s visual features are more discriminative and better suited for retrieval tasks. Between the two DINO variants (base and large), the larger version delivers slightly better performance, the difference is marginal indicating that any DINOv2 backbone can be deployed.
Yo’ LLaVA Dataset
Method/Metric Precision Recall
Positive Negative Weighted
DINOv2 (Large) 74.8 91 98.7 94.9
DINOv2 (Base) 75.5 90.8 98.7 94.7
CLIP (Large) 69 78.3 98.2 88.3
Table 4: Feature Extractor Ablation. We observe that DINO features significantly outperform CLIP features.
A.2 Number of Reference Images (N)
Since our approach does not require a training phase, a key question is how many reference images are needed for robust visual recognition of personalized objects. Figure 4 shows that our method performs well with just one reference image and matches state-of-the-art performance with two images on MyVLM dataset. On Yo’LLaVa dataset, we achieve comparable performance to Yo’LLaVa [20] with only three images, even though the full set includes up to 10 images for some objects.
Figure 4: Ablation on N: Average weighted visual recognition accuracy as a function of number of reference images. Increasing the number of reference images improves performance, but PeKit is robust with just one reference image.
A.3 Number of Personalized Objects (P)
Another key consideration for any personalization method is its robustness to increasing number of personalized objects, particularly as more intra-category instances are introduced into the dataset. Figure 5 illustrates the performance of PeKit on the first 10 objects from the Yo’LLaVA dataset as the number of personalized objects increases incrementally from 10 to all 40 categories. While there is a slight performance drop at higher values of P, PeKit demonstrates overall stability and robustness.
Figure 5: Ablation on P: PeKit’s performance when evaluated on a subset of 10 categories from Yo’LLaVA dataset while progressively increasing number of objects.
A.4 Threshold Selection
In the main paper, we used a fixed threshold of $0.75$ for all settings in order to decide whether a personalized object is present in an image or not. For completeness and flexibility, here we offer a straightforward method for tuning the thresholds on a per-object basis. To adjust the threshold for a given personalized object $i$, we first compute the distribution of distances for the reference images of the object, and then calculate the mean $\mu_{i}$ and standard deviation $\sigma_{i}$ of the distribution. The threshold is then set as $\tau_{i}=\mu_{i}-\sigma_{i}/2$. (We note that by using this method, one can easily adjust the thresholds to be stricter or more loose by adding a scalar multiplier, as $\tau_{i}=\mu_{i}-\alpha(\sigma_{i}/2)$.)
Yo’ LLaVA Dataset
Method/Metric Precision Recall
Positive Negative Weighted
Fixed $\tau=0.75$ 74.8 91 98.7 94.9
Tuned $\tau=\mu-\sigma/2$ 66 92.5 98.1 95.3
Table 5: Automatic object threshold selection.
A.5 Semantic Category
As mentioned in Section 3.2 of the main paper, we employ each object’s semantic category $k_{p}$ on the reference image $I_{p}$ to extract its object-level mask $S_{p}$. For the MyVLM dataset, where the object names are generic, we directly input them into our open-vocabulary detector. Similarly, on This-Is-My-Img dataset, aside from the people in the dataset (queried with the category ‘Man/Woman’), all other object names in this dataset include their semantic category (e.g., Alex’s hat, See Appendix.B).
For the Yo’LLaVA dataset, some concept names such as Vietnamese individual names do not directly indicate the semantic categories. We compare our method’s performance on the Yo’LLaVA dataset by providing the open-vocabulary segmentation model with semantic categories, dataset-provided names, or the generic term ‘main’ during reference view extraction. As shown in Table 6, PeKit achieves state-of-the-art performance even without relying on semantic categories of personalized objects. Note that during inference we consistently use the term ’object’ to extract object proposals using G-Dino.
Method/Metric Precision Positive Acc Negative Acc Weighted Acc
Yo’ LLaVA - 94.9 89.8 92.4
PeKit (categories) 74.8 91 98.7 94.9
PeKit (‘main’) 72.9 86.9 98.9 92.9
PeKit (Vietnamese names) 73.8 88 98.9 93.45
Table 6: Reference View Vocabulary Ablation on Yo’LLaVA Dataset. Our method achieves SOTA results even using the category ’main’ to extract the personalized objects from the reference views.
A.6 Video-QA Object Naming
Object names often offer important contextual clues about the subject in a query to the base LVLM. For example, when asked ”What is Jack doing?”, the model may infer that ”Jack” refers to a male individual in the image. In such cases, the LVLM can often respond accurately without the need for personalization. To isolate the effect of personalization, we perform an ablation study on an open-ended video QA task, where all identified personalized objects are visually prompted using the generic label ENTITY (see Appendix C and Fig.7). Table 7 compares our method’s perfromance when using object names or the term ENTITY to the base LVLM models.
Model/Variation Base PeKit Base+ENTITY PeKit +ENTITY
LLaVA-OneVision 23.0 35.0 14.0 19.9
InternVideoChat 56.6 61.3 45.8 56.2
Table 7: Example table showing values for different VLM model variations
A.7 Time and Memory Requirements
Table 8 details the memory usage and runtime of the various modules in our method for the LVLM personalization task. By default, we employ GroundingDINO (Base) [18] and SAM (Base) [11] for extracting views from reference images and proposing objects during inference, DINOv2 (Large) features for retrieving personalized instances, and LLaVA 1.5 13B as our LVLM for generating answers. All experiments were conducted using an A5000 RTX GPU. As shown in the table, our plug-and-play modules are efficient, adding minimal overhead to the original LLaVA model.
Specifically, our approach introduces an overhead of 0.35 seconds per image (0.31 seconds from G-DINO) relative to LVLM’s inference time. This overhead can be substantially reduced by employing a lightweight object proposal network such as YOLO-World [5], which runs at 55 FPS compared to G-DINO’s 3 FPS. However, as reducing computational time is not the primary focus of this work, we leave further optimization in this direction for future research.
Module Backbone Time/Image (S) Memory (GB)
View Extraction G-DINO + SAM 0.48 1.2
Object Proposal G-DINO 0.31 0.87
Retrieval DINOv2 0.04 1.8
Reasoning LLaVA 1.5 13B 0.87 16.8
Table 8: Time and Memory Requirement. View extraction is conducted only on the reference views corresponding to each personalized object category, whereas the remaining modules are executed for every image during inference.
Appendix B This-Is-My-Img Benchmark Details
The original This-Is-My dataset [30], designed for video-level detection of personalized objects, comprises 104 training and 582 validation short video segments spanning 15 categories. However, some of the segments are no longer available for download. Consequently, one personalized object category, ‘Alex’s Piano’ has been removed from our proposed benchmark due to the unavailability of its training segments.
B.1 Single-concept Set
Table 9 presents further details about the personalized objects and the number of frames included in the single-concept validation splits of our benchmark. We sampled every 10th frame of each validation video segment and manually assigned frames to the respective splits. Note that three categories lack a Negative (Hard) set because the corresponding objects are present in every frame of their validation segments. The number of Positive, Negative (Hard), and Negative (Other) frames varies between categories due to differences in segment length and the number of segments per object. In contrast, the Fake set was standardized, with 10 validation frames generated for each class. Figure 6 offers additional visual examples from our benchmark.
The single-concept validation also includes an open-ended VQA set comprising 70 images (5 per category). Challenging frames were manually selected, and initial question–answer pairs were generated using the GPT-4O model. These pairs were then manually refined to create a high-quality evaluation set. In line with the Yo’LLaVA framework, answers for this set are presented in an A/B multiple-choice format.
Figure 6: This-Is-My-Img Single-concept Benchmark. Our benchmark includes a wide range of concepts presented in realistic indoor and outdoor environments. Reference views can occasionally be sub-optimal, which increases the difficulty of the task. The positive validation set may contain false positives from within the same semantic category, allowing us to assess a model’s robustness to contextual similarities. The negative (hard) and Negative (fake) validation sets are crafted to challenge model resilience by imitating the appearance of personalized objects or their surrounding context. Additionally, the negative ( ’other’) set—though not shown in this image—includes images of all other personalized objects for each concept, serving to evaluate the method’s robustness to dataset bias. Faces are blurred to ensure compliance with GDPR.
Category Positive Negative (Hard) Negative (Other)
Alex’s Bag 161 43 2096
Alex’s Hat 120 0 2137
Blippi’s Shoes 339 79 1918
Casey’s Boosted Board 35 118 $2$
Casey’s Friend Marlan 24 4 2233
Casey’s Son 46 15 2211
Gab’s Puppy Lili 56 12 2201
Nikki’s Camper Bag 229 76 2028
Nikki’s Car 651 184 1606
Reynard’s Keyboard 162 29 2095
Reynard’s Work Chair 188 59 2069
Sherry’s Road Bike 95 12 2162
Zak’s Dog Coffee 26 0 2231
Zak’s Dog Kona 125 0 2132
Sum 2257 631 29341
Table 9: Single-concept Categories and Number of Frames of This-Is-My-Img Benchmark. The validation frames in the dataset are divided into two sets for each category: one with the object visible (Hard) in the image and one without (Other). The benchmark also includes a GPT-genrated set o 10 images per category (Fake).
B.2 Multi-concept Set
Table 10 outlines the category pairs used in our multi-concept validation benchmark. This benchmark comprises 55 images, with each category pair represented by 5 manually selected positive examples sourced from the validation frames. We extended the original This-Is-My-Img dataset categories (Table 9) by incorporating new personalized categories: Alex, Blippi, Casey, Gab, Nikki, Sherry, and Zak. For each of these added categories, we selected 5 reference views from the training segments. Each image in the benchmark is accompanied by an open-ended question–answer pair, collaboratively crafted with GPT-4O. This open-ended visual question answering (VQA) format raises the task’s difficulty and reduces the likelihood that the underlying LVLM model can succeed through answer-choice elimination alone.
Category-pair
Alex - Alex’s Bag Nikki - Nikki’s Car
Alex - Alex’s Hat Nikki - Nikki’s Camper Bag
Blippi - Blippi’s Shoes Sherry - Sherry’s Road Bike
Casey - Casey’s Boosted Board Zak - Zak’s Dog Coffee
Casey - Casey’s Son Zak - Zak’s Dog Kona
Gab - Gab’s Puppy Lili
Table 10: Multi-concept Categories on This-Is-My-Img benchmark. Each category-pair comes with 5 positive frames and 50 negative frames from other category pairs.
B.3 Video-QA Set
As shown in the main paper, PeKit can be extended to video personalization by applying the method to sampled frames. To evaluate this, we used validation video segments (typically under 15 seconds) and curated a VQA dataset with 267 high-quality question–answer pairs for the single-concept categories in Table 9. For each video segment, we employed the LLaVA-OneVision-Qwen2-7B model with a 1-in-16 frame sampling rate to generate an initial set of 1,380 question–answer candidates. These were then filtered for duplicates using GPT-3.5 Turbo, reducing the set to 618 pairs. Finally, we manually refined the results to produce a curated set of 267 high-quality QA pairs.
Appendix C Prompt Templates
In this section, we present the specific query format used across different parts of our experiments.
C.1 Visual prompting
As mentioned in the main paper (Sec 3.4), we employ visual prompting to indicate the personalized object’s location in the image. Next we query our LVLM to personalize its answer given the instance’s name and context. Our prompt to the LVLM for various tasks follows this general structure:
In this image (video), the entity enclosed in a ‘COLOR’ box is called ‘NAME’.
Without mentioning the bounding box and its color, ‘TASK’.
[Optional] Give more details using the information from ‘CONTEXT’.
The COLOR placeholder indicates the color of the bounding box overlaid on the image, the NAME placeholder specifies the instance name, and the TASK placeholder contains the task-specific query. Optionally, the CONTEXT placeholder can contain the prior knowledge about the personalized object retrieved from the memory module.
Note that when multiple objects are detected in one image, the query’s grammatical structure changes to a plural format, and the COLOR and NAME placeholders will contain multiple values separated by commas. The CONTEXT placeholder for each object is included in angle brackets (<>absent<>< >) and contains the NAME for the corresponding instance.
For the experiments in the paper the TASK placeholder can be any of the following prompts:
Personalized Captioning: Describe what is ‘NAME’ doing. Describe the image too.
VQA: Answer the following question about ‘NAME’: ‘QUESTION’.
Figure 7 illustrates an example of our full prompt for each one of the tasks.
Figure 7: Prompt Format. Personalized VQA and captioning on Yo’LLaVA (Left) and MyVLM (Right) datasets. The context used for the ‘red chicken’ is imaginary and generated by ChatGPT.
C.2 Open-ended VQA Validation
To evaluate the accuracy of model predictions in an open-ended VQA task, we adopt the evaluation pipeline proposed by [19]. In particular, we utilize the following prompt template to query a GPT-3.5-Turbo model for assessing the semantic alignment between predicted answers and ground truth responses for each question in the VQA dataset.
‘You are an intelligent chatbot designed for evaluating the correctness of generative outputs for question-answer pairs. Your task is to compare the predicted answer with the correct answer and determine if they match meaningfully.
Here’s how you can accomplish the task: INSTRUCTIONS:
  • Focus on the meaningful match between the predicted answer and the correct answer.
  • Consider synonyms or paraphrases as valid matches.
  • Evaluate the correctness of the prediction compared to the answer.
Please evaluate the following video-based question-answer pair:
Question: QUESTION
Correct Answer: ANSWER
Predicted Answer: PREDICTION
Provide your evaluation only as a yes/no answer. Please generate the response in the form of a Python dictionary string with key ‘pred’, where value of ‘pred’ is a string of ‘yes’ or ‘no’.
DO NOT PROVIDE ANY OTHER OUTPUT TEXT OR EXPLANATION. Only provide the Python dictionary string. For example, your response should look like this: ‘pred’: ‘yes’.’
We calculate the VQA accuracy by dividing the number of times the GPT model, using the specified prompt template, responds with ‘yes’ by the total number of question-answer pairs in the set.
Appendix D Limitations
In this section we discuss limitations of our method and the room for future research.
D.1 Wrong Reference Views.
Our approach depends on extracting instance masks through a segmentation network. Noisy or incorrect masks can increase the likelihood of false positives during inference. As shown in Figure 8, noisy masks, overlaid with red on the personalized object, can lead to incorrect matching during inference, depicted with red bounding boxes. Incorporating a clustering method on the extracted reference features to remove the wrong views might enhance the performance of our method by mitigating this issue.
Figure 8: Noisy Reference Views: Poor segmentation masks may affect the visual prompting stage and degrade PeKit’s performance.
Figure 9: Small Reference Objects: The native image resolution ($518\times 518$) and stride factor (14) of DinoV2 might result in embeddings of small personalized objects, such as Blippi’s shoes, capturing only general attributes, which can increase the likelihood of false positive detections. The incorrect detections are depicted on our proposed Fake validation set. Faces are blurred to ensure compliance with GDPR.
D.2 Small Objects.
Our method performs feature matching at a resolution lower than the original image due to the stride factor (=14) in the DINOv2 encoder. This may lead to suboptimal performance for personalized objects that appear small in the reference images. As shown in Figure 9, with a limited number of patches for the personalized object, DinoV2 may not encode enough fine-grained details, capturing only general attributes like the semantic category (e.g., sneakers) or color, which can result in a higher false positive rate. In case the original reference images are available in high resolution, a potential solution is to crop the (small) personalized objects and resize them to DinoV2’s native resolution ($518\times 518$) before extracting the embeddings.
Figure 10: Comparison to LLaVA: Right: Our method detects personalized objects and integrates provided context (for qualitative comparison) in caption generation. Left: While the original model struggles with specific questions about named objects, our method easily identifies the referred object. Faces are blurred to ensure compliance with GDPR.
D.3 Object Correlations.
Our method, which relies on instance-level detection of personalized objects, could benefit from incorporating contextual information around these objects. This is particularly relevant in scenarios with look-alike objects in contexts that differ from the reference images. For instance, Alex’s everyday bag is more likely to appear in an image featuring Alex rather than someone else. As a potential future direction, we plan to explore methods for extracting object relationships from reference images and integrating them into our LVLM as prior knowledge.
Appendix E Qualitative Comparison to LLaVA
Figure 10 presents examples of PeKit model compared to the base LLaVA [17] on VQA and personalized captioning tasks using images from all three benchmarks discussed in the main paper. When LLaVA does not recognize an object from the given name in the query, it makes guesses, leading to hallucinations or incorrect statements. In contrast, PeKit accurately identifies objects and uses in-context information to guide LLaVA in answering questions or providing details about the image, effectively incorporating in-context information and object appearance. While more advanced prompting or personalized response examples could enhance PeKit, we opted for simplicity and standard design, leaving such improvements for future work.
Appendix F Qualitative Comparison to MyVLM [2]
Figure 11 compares PeKit with MyVLM on images from the MyVLM dataset. We used the original checkpoints and code provided by the authors of MyVLM to generate the results. Checkpoints for the VQA task were not provided.
MyVLM shares the same limitation as Yo’LLaVA, functioning exclusively when the concept identifier is included in the query. Additionally, MyVLM exhibits low precision in detecting personalized objects, as shown in the main paper for the This-Is-My-Img benchmark. This can lead to misidentifying objects as personalized ones. The leftmost image in Figure 11 demonstrates this limitation. MyVLM is asked to provide a caption for the target concept ‘Cat Statue’ while the provided image includes another personalized object, ‘Asian Doll.’ As shown, MyVLM incorrectly identifies the ‘Asian Doll’ as the ‘Cat Statue’ and generates an incorrect personalized caption. Our method addresses this issue by first detecting the correct personalized object(s) and then generating a caption based on the prompt template provided in Appendix C.1.
Furthermore, MyVLM’s training appears to degrade the original LVLM’s captioning capabilities, leading to short captions with hallucinations and sometimes incomprehensible text, leading to a low CLIPScore as demonstrated in the main paper.
Figure 11: Comparison to MyVLM: MyVLM often misidentifies personalized objects because of its low precision. In the leftmost figure, when prompted to caption an image containing a ‘Cat Statue’—which is actually absent—MyVLM incorrectly labels the ‘Asian doll’ and the headset as the ‘Cat Statue’ instead of rejecting the query. Additionally, MyVLM training interferes with the original captioning capabilities of the LVLM, leading to hallucinations, short captions, and sometimes incomprehensible text. For each image, ‘Query’ depicts MyVLM’s sytem prompt where the concept identifier $<$sks$>$ is replaced with the personalized object’s name. PeKit employs its own prompt template described in Appendix C.1.
Figure 12: Qualitative Comparison to Yo’LLaVA: Yo’LLaVA’s prompt template requires specifying the personalized object’s identifier in the query (first row), limiting generalization since users must already know which objects are in the image. Using image-level embeddings can also cause confusion between similar objects (e.g., Alex vs. Alex’s bag). Adjusting the LLM’s head weights further harms captioning quality. PeKit achieves better captioning quality without any training. In each example, the ‘Query’ shows Yo’LLaVA’s prompt with the concept identifier replaced by the object’s name and an added system prompt. PeKit uses a different template, detailed in the Appendix C1. Faces are blurred to ensure compliance with GDPR.
Appendix G Qualitative Comparison to Yo’LLaVA [20]
Figure 12 compares PeKit to Yo’LLaVA for VQA and personalized captioning tasks on This-Is-My-Img benchmark. As seen on the first row, Yo’LLaVA’s prompt template requires the query to include the target object’s concept identifier, making it unsuitable for general captioning tasks and tailored for Visual Question Answering. Besides, since Yo’LLaVA operates on image-level embeddings of reference views, it needs clutter-free object-centered reference views of the personalized objects. As seen on the second row, performance can decline if the personalized object is not in the foreground or if there are other objects/people interacting with the personalized object in the reference views. Besides, fine-tuning the last layer of the language model reduces the LVLM’s captioning capabilities for Yo’LLaVA.